# Cover Learning for Large-Scale Topology Representation

### v1

@@ 1 Introduction
Simplicial complexes (Section A.1) are a specific type of hypergraph, and were originally used in topology to model topological spaces combinatorially; see Figure 1 for an example. Simplicial complexes are gaining increasing attention from applied communities due to their usefulness in topological and geometric inference (Oudot, 2015; Boissonnat et al., 2018; Gardner et al., 2022) and in modelling higher order phenomena (Salnikov et al., 2018; Roddenberry et al., 2021; Chen et al., 2022; Maggs et al., 2024).
@@
Learning simplicial complexes. A simplicial complex on a set $X$ consists of a simplicial complex $K$ together with a function $K_{0}\to\mathsf{Parts}(X)$ from the vertices of $K$ to the set of subsets of $X$, the idea being that vertices represent groups of data points, and higher-dimensional simplices represent relationships between these. Simplicial complex learning is the process of constructing a simplicial complex on an input dataset, and it encompasses standard constructions such as graphs in classical unsupervised learning (e.g., neighborhood graphs in Laplacian eigenmaps (Belkin & Niyogi, 2003), t-SNE (Van der Maaten & Hinton, 2008), and DBSCAN (Ester et al., 1996)), simplicial complexes in topological inference (e.g., geometric complexes (Edelsbrunner & Harer, 2022) and Mapper (Singh et al., 2007)), and meshes is surface reconstruction (Dey, 2006).
Learning covers. A cover of a set $X$ is a family $\{U_{i}\}_{i\in I}$ of subsets of $X$, which cover it in the sense that $X=\bigcup_{i\in I}U_{i}$. Cover learning is the process of constructing a cover of an input dataset.
@@
Contributions. We now outline our main contributions.
@@
Structure of the paper. In Section 2, we overview related work, specifically algorithms in TDA that fit the framework of cover learning. In Section 3, we develop theory for cover learning based on optimization, which we use in Section 4 to design a cover learning algorithm. Section 5 contains computational examples showcasing our methods. Appendix A has background, and other appendices contain details about theory and experiments.
@@ 2 Related work: cover learning in TDA
Although the concept of cover learning algorithm is introduced in this paper (Section 1), several unsupervised learning algorithms in TDA are in essence cover learning algorithms. Here, we overview some of these.
Cover learning algorithms with a filter function. To the best of our knowledge, 1D Mapper (Singh et al., 2007) is the first cover learning algorithm designed with the nerve construction in mind (see Algorithm 1 for the cover learning algorithm underlying 1D Mapper). The advantages of 1D Mapper are its simplicity and flexibility: different choices of the function $f:X\to\mathbb{R}$, known as the filter function, can provide widely different outputs, reflecting the structure of the data as seen by the filter function. We now describe 1D Mapper’s main limitations:
@@
    Dependence and choice of cover of range of filter. The initial cover $\{I_{i}\}_{i=1}^{k}$ of $\mathbb{R}$ (or, equivalently, of the range $[\min f,\max f]$ of the filter function) is also difficult to choose. This requires choosing the number of intervals $k$, as well as their endpoints.
@@
    Unsuitability for higher-dimensional topological inference. The nerve of the cover produced by 1D Mapper is $1$-dimensional, and does not contain higher-dimensional information. We elaborate on this in Appendix D.
@@
Recently, and closest to this paper, Differentiable (1D) Mapper (Oulhaj et al., 2024) was introduced to address (1) by using optimization to automatically search for a filter function; this comes at the expense of having to choose a loss function and a parametric family of filter functions.
@@
Algorithm 1 1D Mapper cover learning algorithm
  Input: Data $X$, function $f:X\to\mathbb{R}$, clustering algorithm $C_{\theta}$, parameter(s) $\theta$ for $C_{\theta}$, cover $\{I_{i}\}_{i=1}^{k}$ of $\mathbb{R}$
[…]
  Let $\mathcal{U}_{i}\coloneqq C_{\theta}(f^{-1}(I_{i}))$ for $1\leq i\leq k$
  Return The union $\bigcup_{i=1}^{k}\mathcal{U}_{i}$
Algorithm 2 Ball Mapper cover learning algorithm
  Input: Data $X$, $\varepsilon>0$
  Build an $\varepsilon$-net $\{y_{i}\}_{i=1}^{k}$ of $X$
  Return The cover $\{B(y_{i},\varepsilon)\}_{i=1}^{k}$
Cover learning without a filter function. The category of cover learning algorithms which do not rely on a filter function (implicitly or explicitly) is remarkably small. To the best of our knowledge, there is a single TDA cover learning algorithm in this category: Ball Mapper (Dłotko, 2019; Dlotko, 2019), described in Algorithm 2.
Ball Mapper’s main limitation is its reliance on a global distance scale $\varepsilon>0$, which must be chosen by the user, does not allow for variations in density, and is susceptible to the curse of dimensionality. Of note is the fact that Ball Mapper is essentially equivalent to the Witness complex (De Silva & Carlsson, 2004) with parameter $v=0$ (Appendix E). Because of this, consistency results from TDA can be used to prove consistency results for Ball Mapper, meaning that, contrary to the cover learning algorithms based on 1D filter functions, Ball Mapper can be used to do higher-dimensional topological inference.
@@ 3 Theory
A formalized goal. Condition $(\mathsf{M})$ in 1 is formalized by requiring the cover elements to have small measure, condition $(\mathsf{G})$ by asking that the cover elements have boundary of small measure, while condition $(\mathsf{T})$ is formalized by asking the cover to be homology-good, which ensures that the nerve recovers the correct homology, by the nerve theorem.
@@
Fuzzy covers. Observe that a $k$-element cover of a set, i.e., a cover of the form $\{U_{i}\}_{i=1}^{k}$, is the same as a function $g:X\to\{0,1\}^{k}$ such that $\max_{1\leq i\leq k}g_{i}(x)=1$ for all $x\in X$. Here, the $\max$ operation is ensuring that every element belongs to some cover element. Indeed, the equivalence is given by mapping a cover $\{U_{i}\}_{1\leq i\leq k}$ to the vector $(\chi_{U_{1}},\dots,\chi_{U_{k}}):X\to\{0,1\}^{k}$ of indicator functions.
In order to turn the space of covers into a continuous space, we first replace the two-element set $\{0,1\}$ by the unit interval $[0,1]$, and define
@@
In order to produce a single cover out of a fuzzy cover, one can threshold it, as follows. Given any function $f:X\to[0,1]$ and a threshold $\lambda\in[0,1]$, consider the suplevel set $\left\{f>\lambda\right\}\coloneqq\{x\in X:f(x)>\lambda\}$. Then, given a fuzzy cover $g:X\to\Gamma^{k-1}$ and a $\lambda\in[0,1]$, we can threshold the fuzzy cover $g$ to obtain a cover of $X$, as follows:
@@
Although this will be of use later, let us remark here that a fuzzy cover $g:X\to\Gamma^{k-1}$ gives rise to a filtered simplicial complex $\{\mathsf{Ner}(\left\{g>\lambda\right\})\}_{\lambda\in[0,1]}$; see Appendix C.
Random covers from fuzzy covers. Suppose that $g:X\to\Gamma^{k-1}$ is a fuzzy cover of $X$. By sampling a threshold $\lambda\in[0,1]$ uniformly at random, and considering $\left\{g>\lambda\right\}$, we get a random cover of $X$, so each fuzzy cover $g$ induces a probability measure $\mu_{g}$ on the space of covers of $X$. Formally, $\mu_{g}$ is the pushforward of the uniform measure on $[0,1]$ along the function $\lambda\mapsto\left\{g>\lambda\right\}$ (see Section B.3 for details).
Random covers in the context of TDA were first considered in (Oulhaj et al., 2024), with the similar goal of producing a cover of a dataset via optimization (Section 2). Their methodology, however, follows the Mapper pipeline and relies on filter functions.
The idealized loss function. We now describe our idealized loss function for fuzzy cover learning, which is essentially obtained by taking the expected value of the conditions in 2 for a random cover distributed according to the probability measure $\mu_{g}$ corresponding to a fuzzy cover $g$.
@@
[…]
We now state our main theoretical result, which says that the terms in 1 can be either computed or bounded from above explicitly using the fuzzy cover $g$.
@@ 4 Practice
Our practical approach to cover learning relies on two main considerations: estimating the loss of 1 (Section 4.1), and finding suitable parametrizations of the space of fuzzy covers (Section 4.2). Other practical considerations are required to make optimization feasible and efficient (Section 4.3). Building on these, we describe the ShapeDiscover cover learning algorithm in Section 4.4.
@@ 4.1 Loss estimation
Let $X$ be the input data (seen as a set), and let $G$ be a weighted graph, with $X$ as its set of vertices, and with weight $w(x,y)$ for $(x,y)\in G$ representing relationship strength between $x$ and $y$. Let $g:X\to[0,1]^{k}$ (for example, $g$ could be a fuzzy cover). Based on Theorem 3.2, we define the following estimators for the losses:
@@
where $\eta_{M},\eta_{G},\eta_{T},\eta_{R}$ are normalization constants (depending only on the graph) for the purpose of having the losses be in similar ranges; see Section F.1.
For the topological loss, we use the relaxation which only considers 00-dimensional homology (Section B.2), since 00-dimensional homology is more efficient to compute than higher-dimensional homology, and still performs well in the topological inference examples (Section 5.1).
@@ 4.2 Parametric fuzzy covers
In order to make our approach flexible enough to be combined with standard machine learning models, we first describe a way of turning any parametric vector-valued model $f_{\theta}:X\to\mathbb{R}^{k}$ on the input data $X$ into one of the form $g_{\theta}:X\to\Gamma^{k-1}$, thus parametrizing the space of fuzzy covers on the data. See the end of this section for possible initial models $f_{\theta}$, including neural networks and graph neural networks.
@@
so that, in particular, $\Delta^{k-1}_{\infty}=\Gamma^{k-1}$, and $\Delta^{k-1}_{1}$ is the standard, $(k-1)$-dimensional simplex. A function $h:X\to\Delta_{1}^{k-1}$ is called a partition of unity, since it consists of functions $h_{1},\dots,h_{k}:X\to[0,1]$ such that $\sum_{i}h_{i}\equiv 1$. Note that, by definition of the standard softmax function, we have $\mathsf{softmax}:\mathbb{R}^{k}\to\Delta_{1}^{k-1}$. Thus, given any parametric model $f_{\theta}:X\to\mathbb{R}^{k}$, we obtain a parametric partition of unity by simply postcomposing with $\mathsf{softmax}$.
@@
Possible parametric models. Depending on the type of data $X=\{x_{1},\dots,x_{n}\}$, there are different choices for $f_{\theta}$:
  • The simplest option to let $\theta$ be an $n$-by-$k$ matrix, and to let $f_{\theta}(x_{i})_{j}=\theta_{i,j}$.
@@ 4.3 Other implementation details
Initialization. In order to make optimization fast and reliable, it is convenient to use a good initialization. Any clustering of a set is in particular a cover, and we propose to use a geometrically meaningful clustering (obtained by, e.g., $k$-means or spectral clustering) as initial fuzzy cover.
Smooth parametric fuzzy covers. As described in Section 4.2, if $f_{\theta}:X\to\mathbb{R}^{k}$ is a parametric vector-valued model, then $\pi_{\infty}\circ\mathsf{softmax}\circ f_{\theta}:X\to\Delta^{k-1}_{\infty}=\Gamma^{k-1}$ is a parametric fuzzy cover. However, the operation $\pi_{\infty}$ given by normalization by the $\infty$-norm is not smooth; we thus propose to replace $\pi_{\infty}$ with $\pi_{p}$ for some $p\in[1,\infty)$ during optimization, and to only use $\pi_{\infty}$ to obtain the final fuzzy cover; see Algorithm 3. We expect larger values of $k$ to require larger values of $p$, but this approximation does not seem to influence results significantly, and we let $p=5$ in all experiments.
Topological optimization. In order to optimize the loss $\widehat{\mathsf{T}}_{0}$ we use the topological optimization with big steps of (Nigmetov & Morozov, 2024); see Section A.3.4.

### v2

@@ 1.1 Context
Simplicial complexes (Section A.1) are a specific type of hypergraph, and were originally used in topology to model topological spaces combinatorially; see Figure 1 for an example. Recent applications include neuroscience (Gardner et al., 2022), biology (Benjamin et al., 2024), physics (Sale et al., 2023), and machine learning (Maggs et al., 2024).
@@
1.2 Learning simplicial complexes
A simplicial complex on a set $X$ consists of a simplicial complex $K$ together with a function $K_{0}\to\mathsf{Parts}(X)$ from the vertices of $K$ to the set of subsets of $X$, the idea being that vertices represent groups of data points, and higher-dimensional simplices represent relationships between these. Simplicial complex learning is the process of constructing a simplicial complex on an input dataset, and it encompasses standard constructions such as graphs in classical unsupervised learning (e.g., neighborhood graphs in Laplacian eigenmaps (Belkin & Niyogi, 2003), t-SNE (Van der Maaten & Hinton, 2008), and DBSCAN (Ester et al., 1996)), simplicial complexes in topological inference (e.g., geometric complexes (Edelsbrunner & Harer, 2022) and Mapper (Singh et al., 2007)), and meshes is surface reconstruction (Dey, 2006).
1.3 Learning covers
A cover of a set $X$ is a family $\{U_{i}\}_{i\in I}$ of subsets of $X$, which cover it in the sense that $X=\bigcup_{i\in I}U_{i}$. Cover learning is the process of constructing a cover of an input dataset.
@@ 1.3 Learning covers
1.4 Cover learning vs. clustering
The clustering problem can be understood as the zero-dimensional case of the cover learning problem, in which the relevant geometry of the data is entirely determined by the connected components, with no relevant topological information above dimension zero. Although the cover learning problem cannot be completely reduced to the clustering problem, our proposed method does rely on clustering for initialization; see Section 4.3.
1.5 Contributions
We describe our four main contributions.
@@ 1.5 Contributions
1.6 Structure of the paper
In Section 2, we overview related work, specifically algorithms in TDA that fit the framework of cover learning. In Section 3, we develop theory for cover learning based on optimization, which we use in Section 4 to design a cover learning algorithm. Section 5 contains computational examples showcasing our methods. Appendix A has background, and other appendices contain details about theory and experiments.
@@ 2 Related work: cover learning in TDA
Although the concept of cover learning algorithm is introduced in this paper (Section 1), several unsupervised learning algorithms in TDA are in essence cover learning algorithms. Here, we overview some of these. For other related work, see Appendix E.
Cover learning algorithms with a filter function. To the best of our knowledge, 1D Mapper (Singh et al., 2007) is the first cover learning algorithm designed with the nerve construction in mind (see Algorithm 1 for the cover learning algorithm underlying 1D Mapper). The advantages of 1D Mapper are its simplicity and flexibility: different choices of the function $f:X\to\mathbb{R}$, known as the filter function, can provide widely different outputs, reflecting the structure of the data as seen by the filter function. We now describe four of 1D Mapper’s main limitations:
@@
    Dependence and choice of cover of range of filter. The initial cover $\{I_{i}\}_{i=1}^{k}$ of $\mathbb{R}$ (or, equivalently, of the range $\min f,\max f]$ of the filter function) is also difficult to choose. This requires choosing the number of intervals $k$, as well as their endpoints.
@@
    Unsuitability for higher-dimensional topological inference. The nerve of the cover produced by 1D Mapper is $1$-dimensional, and does not contain higher-dimensional information. We elaborate on this in Appendix F.
Algorithm 1 1D Mapper cover learning algorithm
Input: Data $X$, function $f:X\to\mathbb{R}$, clustering algorithm $C_{\theta}$, parameter(s) $\theta$ for $C_{\theta}$, cover $\{I_{i}\}_{i=1}^{k}$ of $\mathbb{R}$.Take pullback cover $\{f^{-1}(I_{i})\}_{i=1}^{k}$ of $X$
 Let $\mathcal{U}_{i}\coloneqq C_{\theta}(f^{-1}(I_{i}))$ for $1\leq i\leq k$
Return The union $\bigcup_{i=1}^{k}\mathcal{U}_{i}$.
Algorithm 2 Ball Mapper cover learning algorithm
Input: Data $X$, $\varepsilon>0$.
 Build an $\varepsilon$-net $\{y_{i}\}_{i=1}^{k}$ of $X$
Return The cover $\{B(y_{i},\varepsilon)\}_{i=1}^{k}$.
@@
Recently, and closest to this paper, Differentiable 1D Mapper (Oulhaj et al., 2024) was introduced to address (1) by using optimization to automatically search for a filter function; this comes at the expense of having to choose a loss function and a parametric family of filter functions.
@@
Cover learning without a filter function. The category of cover learning algorithms which do not rely on a filter function (implicitly or explicitly) is remarkably small. To the best of our knowledge, there is a single TDA cover learning algorithm in this category: Ball Mapper (Dłotko, 2019a, b), described in Algorithm 2. Consistency results can be proven for Ball Mapper, which imply that, contrary to cover learning algorithms based on 1D filter functions, Ball Mapper can be used to do higher-dimensional topological inference.
Ball Mapper’s main limitation is its reliance on a global distance scale $\varepsilon>0$, which must be chosen by the user, does not allow for variations in density, and is susceptible to the curse of dimensionality. Of note is the fact that Ball Mapper is essentially equivalent to the Witness complex (De Silva & Carlsson, 2004) with parameter $v=0$ (Appendix G).
@@ 3 Theory
3.1 Formal goal
Condition $(\mathsf{M})$ in 1 is formalized by requiring the cover elements to have small measure, condition $(\mathsf{G})$ by asking that the cover elements have boundary of small measure, while condition $(\mathsf{T})$ is formalized by asking the cover to be homology-good, which ensures that the nerve recovers the correct homology, by the nerve theorem.
@@ 3.2 Optimization over spaces of covers
3.2.1 Fuzzy covers
Observe that a $k$-element cover of a set, i.e., a cover of the form $\{U_{i}\}_{i=1}^{k}$, is the same as a function $g:X\to\{0,1\}^{k}$ such that $\max_{1\leq i\leq k}g_{i}(x)=1$ for all $x\in X$. Here, the $\max$ operation is ensuring that every element belongs to some cover element. Indeed, the equivalence is given by mapping a cover $\{U_{i}\}_{1\leq i\leq k}$ to the vector $(\chi_{U_{1}},\dots,\chi_{U_{k}}):X\to\{0,1\}^{k}$ of indicator functions.
In order to turn the space of covers into a continuous space, we replace the two-element set $\{0,1\}$ by the unit interval $[0,1]$, and define
@@ 3.2.1 Fuzzy covers
The exponent k1k-1 in $\Gamma^{k-1}$ represents the topological dimension of the space. This notation is consistent with the usual convention that denotes the standard simplex with $k$ vertices by $\Delta^{k-1}$ (see also Section 4.2.1).
@@
In order to produce a single cover out of a fuzzy cover, one can threshold it, as follows. Given any function $f:X\to[0,1]$ and a threshold $\lambda\in[0,1]$, consider the suplevel set $\left\{f>\lambda\right\}\coloneqq\{x\in X:f(x)>\lambda\}$. Then, given a fuzzy cover $g:X\to\Gamma^{k-1}$ and $\lambda\in[0,1]$, we can threshold the fuzzy cover $g$ to obtain a cover of $X$, as follows:
@@
Since this will be of use later, let us remark here that a fuzzy cover $g:X\to\Gamma^{k-1}$ gives rise to a filtered simplicial complex $\{\mathsf{Ner}(\left\{g>\lambda\right\})\}_{\lambda\in[0,1]}$ (see Appendix D).
3.2.2 Random covers from fuzzy covers
Suppose that $g:X\to\Gamma^{k-1}$ is a fuzzy cover of $X$. By sampling a threshold $\lambda\in[0,1]$ uniformly at random, and considering $\left\{g>\lambda\right\}$, we get a random cover of $X$, so each fuzzy cover $g$ induces a probability measure $\mu_{g}$ on the space of covers of $X$. Formally, the probability measure $\mu_{g}$ is the pushforward of the uniform measure on $[0,1]$ along the function $\lambda\mapsto\left\{g>\lambda\right\}$ (see Section B.3 for details).
Random covers in the context of TDA were first considered in (Oulhaj et al., 2024), with the similar goal of producing a cover of a dataset via optimization (Section 2). Their methodology, however, follows the 1D Mapper pipeline, and relies on filter functions.
3.3 The idealized loss function
Loss function for a Riemannian manifold. Our idealized loss function for fuzzy cover learning is obtained by taking the expected value of the conditions in 2 for a random cover distributed according to the probability measure $\mu_{g}$ corresponding to a fuzzy cover $g$.
@@ 3.3 The idealized loss function
$\mathop{\text{\large min}}\limits_{\begin{subarray}{c}g:\mathcal{M}\to\Gamma^{k-1}\\ \text{smooth}\end{subarray}}\;\;\alpha_{M}\cdot\mathsf{M}(g)\;+\;\alpha_{G}\cdot\mathsf{G}(g)\;+\;\alpha_{T}\cdot\mathsf{T}(g)$ $\mathsf{M}(g)$ $\coloneqq\sum_{i=1}^{k}\;\mathbb{E}\Big(\,\mathcal{H}^{d}\big(U_{i}\big)\,\Big)^{2}$ $\mathsf{G}(g)$ $\coloneqq\sum_{i=1}^{k}\;\mathbb{E}\Big(\,\mathcal{H}^{d-1}\big(\partial U_{i}\big)\,\Big)^{2}$ $\mathsf{T}(g)$ $\coloneqq\mathbb{P}\Big(\left\{U_{i}\right\}_{i=1}^{k}\text{ is not homology-good }\Big)$ $\{U_{i}\}_{i=1}^{k}$ random cover with probability measure $\mu_{g}$
Next is our main theoretical result, which allows us to minimize the loss terms in 1 in practice.
@@ 4 Practice
Our practical approach to cover learning relies on three main considerations: estimating the loss of 1 (Section 4.1), finding a suitable parametrization of the space of fuzzy covers (Section 4.2), and computing a good initialization (Section 4.3). Other practical considerations are required to make optimization feasible and efficient (Section 4.4). Building on these, we describe the ShapeDiscover cover learning algorithm in Section 4.5.
@@ 4.1 Loss estimation
Let $X$ be the input data (seen as a set), and let $G$ be a weighted graph with $X$ as its set of vertices, and with weight $w(x,y)$ for $(x,y)\in G$ representing relationship strength between $x$ and $y$. Let $g:X\to[0,1]^{k}$ (for example, $g$ could be a fuzzy cover). Based on Theorem 3.2, we define the following estimators for the losses:
@@
where $\eta_{M},\eta_{G},\eta_{T},\eta_{R}$ are normalization constants (depending only on the graph) for the purpose of having the losses be in similar ranges; see Section H.1.
For the topological loss, we use the relaxation which only considers 0-dimensional homology (Section B.2), since 0-dimensional homology is more efficient to compute than higher-dimensional homology, and still performs well in the examples (Section 5.1).
@@ 4.2 Parametric fuzzy covers
In order to make our approach flexible enough to be combined with standard machine learning models, we first describe a way of turning any parametric vector-valued model $f_{\theta}:X\to\mathbb{R}^{k}$ on the input data $X$ into one of the form $g_{\theta}:X\to\Gamma^{k-1}$, thus parametrizing the space of fuzzy covers on the data. See Section 4.2.2 for possible initial models $f_{\theta}$, including neural networks and graph neural networks.
4.2.1 Parametrizing fuzzy covers with softmax
@@
so that, in particular, $\Delta^{k-1}_{\infty}=\Gamma^{k-1}$, and $\Delta^{k-1}_{1}$ is the standard (k1)(k-1)-dimensional simplex. A function $h:X\to\Delta_{1}^{k-1}$ is called a partition of unity, since it consists of functions $h_{1},\dots,h_{k}:X\to[0,1]$ such that $\sum_{i}h_{i}\equiv 1$. Note that, by definition of the standard softmax function, we have $\mathsf{softmax}:\mathbb{R}^{k}\to\Delta_{1}^{k-1}$. Thus, given any parametric model $f_{\theta}:X\to\mathbb{R}^{k}$, we obtain a parametric partition of unity by simply postcomposing with $\mathsf{softmax}$.
@@
4.2.2 Parametric models
Depending on the type of data $X=\{x_{1},\dots,x_{n}\}$, there are different choices for $f_{\theta}$:
  • The simplest option to let $\theta$ be an $n$-by-$k$ matrix, and to set $f_{\theta}(x_{i})_{j}=\theta_{i,j}$.
@@ 4.2.2 Parametric models
4.3 Initialization
In order to make optimization fast and reliable, it is necessary to use a good initialization. Any clustering of a set is in particular a cover, and we propose to use a geometrically meaningful clustering (obtained by, e.g., $k$-means or spectral clustering) as initial fuzzy cover.
@@ 4.4 Other implementation details
4.4.1 Smooth parametric fuzzy covers
As described in Section 4.2, if $f_{\theta}:X\to\mathbb{R}^{k}$ is a parametric vector-valued model, then $\pi_{\infty}\circ\mathsf{softmax}\circ f_{\theta}:X\to\Delta^{k-1}_{\infty}=\Gamma^{k-1}$ is a parametric fuzzy cover. However, the operation $\pi_{\infty}$ given by normalization by the $\infty$-norm is not smooth. We thus propose to replace $\pi_{\infty}$ with $\pi_{p}$ for some $p\in[1,\infty)$ during optimization, and to only use $\pi_{\infty}$ to obtain the final fuzzy cover; see Algorithm 3. We expect larger values of $k$ to require larger values of $p$, but this approximation does not seem to influence results significantly, and we let $p=5$ in all experiments.
4.4.2 Topological optimization
In order to optimize the loss $\widehat{\mathsf{T}}_{0}$ we use the topological optimization with big steps of (Nigmetov & Morozov, 2024); see Section A.3.4.